A cleaner pushes a 4.50-kg laundry cart in such a way that the net external force on it is 60.0N. Calculate its acceleration.
1
Expert's answer
2019-07-01T09:26:26-0400

We can find the acceleration of the laundry cart from the Newton's Second Law of Motion:


"F_{net} = ma,"

here, "F_{net} = 60.0 N" is the net external force acting on the cart, "m = 4.50 kg" is the mass of the cart and "a" is the acceleration of the cart.

Then, from this formula we can calculate the acceleration of the laundry cart:


"a = \dfrac{F_{net}}{m} = \dfrac{60.0 N}{4.50 kg} = 13.3 \dfrac{m}{s^2}."

Answer:

"a = 13.3 \dfrac{m}{s^2}."